I just need to know what mathematical equation to use to find out how old I will be when I'm 3x's older than someone? I was 12 when my little sister was born, and I need to find out how old we will each be when I'm 3x's older than her. Thanks for your help.
Answer by sudhanshu_kmr(1152)


let your age =x and sister age = y
difference between ages i.e x - y = 12
( the difference between age of two always remain same)

and according to question, x = 3y

so, 3y -y = 12
=> 2y = 12
=> y = 12/2 = 6
x = 18
your age = 18, sister age = 6